# Calib3D: Calibrating Model Preferences for Reliable 3D Scene Understanding

### v1

@@ Abstract
Safety-critical 3D scene understanding tasks necessitate not only accurate but also confident predictions from 3D perception models. This study introduces Calib3D, a pioneering effort to benchmark and scrutinize the reliability of 3D scene understanding models from an uncertainty estimation viewpoint. We comprehensively evaluate 28 state-of-the-art models across 10 diverse 3D datasets, uncovering insightful phenomena that cope with both the aleatoric and epistemic uncertainties in 3D scene understanding. We discover that despite achieving impressive levels of accuracy, existing models frequently fail to provide reliable uncertainty estimates – a pitfall that critically undermines their applicability in safety-sensitive contexts. Through extensive analysis of key factors such as network capacity, LiDAR representations, rasterization resolutions, and 3D data augmentation techniques, we correlate these aspects directly with the model calibration efficacy. Furthermore, we introduce DeptS, a novel depth-aware scaling approach aimed at enhancing 3D model calibration. Extensive experiments across a wide range of configurations validate the superiority of our method. We hope this work could serve as a cornerstone for fostering reliable 3D scene understanding. Code and benchmark toolkits are publicly available. [^1] GitHub Repo: https://github.com/ldkong1205/Calib3D.
Keywords:
3D Semantic Segmentation Uncertainty Estimation Network Calibration 3D Perception Robustness Autonomous Driving
@@ 1 Introduction
The reliability of the perception system in real-world conditions is paramount. Safety-critical applications, such as autonomous driving and robot navigation, often rely on robust and accurate predictions from perception models [130, 98].
Albeit being the most popular choice, learning-based perception models are known to struggle in delivering uncertainty estimates [34] and suffer from over- or under-confidences [47]. In other words, prior models tend to be badly calibrated, which cannot satisfy the needs in real life [1, 79]. This is contradictory to the objective of safe operations of autonomous vehicles, where the perception models are anticipated to predict the presence or absence of obstacles in a fairly confident manner [97, 55, 116]. A similar rationale often applies to other safety-critical applications, such as surveillance [62, 77, 88], automated health care [74, 50, 73], remote sensing [94, 33], etc., where the trustworthiness plays the critical role.
Figure 1: Well-calibrated 3D scene understanding models are anticipated to deliver low uncertainties when predictions are accurate and high uncertainties when predictions are inaccurate. Existing 3D models [133] (UnCal) and prior calibration methods [36, 71] struggled to provide proper uncertainty estimates. Our proposed depth-aware scaling (DeptS) is capable of outputting accurate estimates, highlighting its potential for real-world usage. The plots shown are the point-wise expected calibration error (ECE) rates. The colormap goes from dark to light denoting low and high error rates, respectively.
@@
Motivation. Despite existing endeavors, the reliability of 3D scene understanding models in delivering uncertainties [^2] In this work, for the sake of clarity, the terms uncertainty and confidence are used interchangeably, i.e., uncertainty $=1~{}-$ confidence. remains unknown. 3D data, such as those acquired by LiDAR sensors and RGB-D cameras, are much sparser and often lack structures than images [11, 2, 114]. Calib3D emerges as a comprehensive study dedicated to scrutinizing and benchmarking the reliability of 3D scene understanding models through the lens of uncertainty estimation. Our exploration delves into both aleatoric and epistemic uncertainties, aiming to unravel the nuanced dynamics of 3D perception models and their calibration needs in real-world, safety-critical settings. The major focuses of this study are:
@@
  • Aleatoric Uncertainty in 3D. We examine how intrinsic factors, such as sensor measurement noises [31] and point cloud density variations [55, 105], contribute to data uncertainty in 3D scene understanding. Such uncertainty cannot be reduced even with more data or improved models, necessitating efforts that effectively interpret and quantify this inherent variability. Calib3D encompasses a total of 10 popular 3D datasets, ranging from diverse sensor acquisitions, annotation protocols, and scene conditions. Our benchmark covers diverse 3D scenarios, including driving, off-road, and indoor scenes, as well as dynamic, synthetic, corrupted, and adverse weather conditions.
@@
  • Epistemic Uncertainty in 3D. Different from the rather unified network structures in 2D [40, 25], 3D scene understanding models encompass a wider array of structures due to the complex nature of 3D data processing. Our investigation in Calib3D extends to the model uncertainty associated with the diverse 3D architectures, highlighting the importance of addressing knowledge gaps in model training and data representation. A total of 28 state-of-the-art models are compared and analyzed, shedding light on the future development of more reliable 3D scene understanding models.
@@
Our investigation into the current landscape of 3D scene understanding models reveals a critical gap: while these models often achieve promising levels of accuracy, their calibration abilities –essential for trust in safety-critical applications – consistently fall short of the mark. This discrepancy, as visualized in Fig. 1, highlights the need for more sophisticated calibration methods that can align model confidence with their predictive accuracy. In-depth analysis of various factors, including network capacity, LiDAR data representations, rasterization resolutions, and 3D data augmentation techniques, has led us to identify key areas where calibration performance can be significantly improved.
@@
To further enhance uncertainty estimation capabilities, we propose a depth-aware scaling method, dubbed DeptS. Our method is motivated by the observation that uncalibrated models tend to have low accuracy in the middle-to-far region of the ego-vehicle, while, in the meantime, posing severe over-confident predictions. This problem, which is directly correlated with 3D scene structural information, inevitably leads to high calibration errors. To tackle this challenge, we design a depth-correlated temperature to dynamically adjust the logits distribution based on the depth information, exhibiting strong generalizability in calibrating 3D perception models. As shown in Fig. 1, DeptS not only demonstrates appealing calibration performance over the uncalibrated model but also outperforms several calibration methods from existing literature [36, 61, 110, 71].
To encapsulate, this work is featured by the following seminar contributions:
  • To the best of our knowledge, Calib3D is the first benchmark targeted at probing the uncertainties of 3D perception models in real-world conditions.
  • Our benchmark encompasses a systematic study of 28 state-of-the-art models across 10 diverse 3D datasets, laying a solid foundation for the future research and development of reliable 3D scene understanding models.
  • We proposed DeptS, a straightforward yet effective depth-aware scaling method that better calibrates the uncertainty estimates for 3D models.
  • Extensive experimental evaluations across a wide range of 3D scenarios demonstrate our advantages, shedding light on a more reliable 3D scene understanding that extends well beyond the current state of the art.
@@ 2 Related Work
3D Scene Understanding. Holistic perception of 3D scenes lays the foundation of many real-world applications [11, 6]. Based on the choice of 3D representations [106], existing approaches can be categorized into range view [75, 131, 16, 118, 57, 54, 122], bird’s eye view [129, 14, 132], sparse voxel [18, 133, 101, 44, 43, 17], and point-based [45, 103, 128, 89] methods. Most recent works seek to combine the advantages of multiple representations together [65, 119, 67] or fuse other modalities (e.g., cameras, Radars, and IMU) with point clouds [134, 49, 68, 85, 120] for more accurate 3D perceptions. Although promising performance has been achieved on popular 3D scene understanding datasets, the reliability of existing 3D perception models in estimating the predictive uncertainty remains unknown.
@@
Uncertainty Estimation. Quantifying uncertainties is crucial in real-world scenarios [4, 41]. For safety-critical applications like 3D scene understanding, it is especially critical to obtain reliable estimates of uncertainties [82]. In general, methods on the uncertainty estimation track can be categorized into various types, i.e., deterministic networks [72, 96], Bayesian methods [23, 104, 10, 42, 30], ensembles [63, 92, 37], and test-time augmentations [70, 5]. Recent studies pay attention to post-hoc approaches for calibrating uncertainties, which align with practical usages [34]. This work follows this line of research and extends efforts to 3D scene understanding, hoping to enlighten future works on this crucial topic.
@@
Network Calibration. As one of the prevailing research topics, numerous calibration methods have been proposed to tackle different tasks, including image classification [36, 127, 61, 71, 80, 84, 60], semantic segmentation [24, 51, 110, 76, 9, 27], object detection [62, 77, 52, 81, 88], depth estimation [53, 48, 87, 108], remote sensing [94, 33], medical imaging [74, 50, 73], and so on. However, the calibration of 3D scene understanding models, which relates closely to real-world applications, is rather overlooked in the literature. Dreissig et al. [28] made a preliminary study of the uncertainties in SalsaNext [21] on SemanticKITTI [7]. As far as we know, Calib3D serves as the first systematic study of uncertainty estimation for 3D scene understanding. Our benchmark covers a total of 28 state-of-the-art models across 10 diverse 3D datasets, paving the way for the future development of more reliable and practical 3D scene understanding models. Additionally, we contribute a novel depth-aware scaling method, DeptS, that can effectively calibrate the uncertainty estimates of 3D scene understanding models.
3D Robustness. Several recent studies have been presented to benchmark the robustness of 3D perception models for driving applications. One line of research aims to probe the model’s 3D robustness against common corruptions from point clouds [55, 39, 38, 93, 12, 124, 115], depth [59, 32, 56], multi-view images [116, 46, 15], etc. Other related works explore the 3D robustness from aspects like sensor failures [126, 35] and adversarial attacks [130, 117]. In this work, different from all previous pursuits, we target another important direction on this relevant topic. As far as we know, Calib3D establishes the first 3D robustness benchmark from the uncertainty estimation perspective. Extensive experiments of 3D scene understanding models on diverse datasets are conducted, which could shed light on the future development of practical and reliable 3D scene understanding models.
@@ 3 Calib3D
A typical point cloud data $\mathcal{P}=\{\mathbf{p}_{i},q_{i}|i=0,1,...,N-1\}$ consists of a total of $N$ points acquired by the sensor, where $\mathbf{p}_{i}\in\mathbb{R}^{3}$ holds the Cartesian coordinates $(p^{x}_{i},p^{y}_{i},p^{z}_{i})$ and $q_{i}\in\mathbb{R}^{1}$ is the sensor reflection value, e.g., the laser intensity. For a learning-based system, there is an associated semantic label $\mathcal{Y}=\{y_{i}|i=0,1,...,N-1\}$ for each point in $\mathcal{P}$, where $y_{i}$ takes values from $S$ pre-defined semantic classes, depending on the specific label mapping. The random variables $\mathcal{P}$ and $\mathcal{Y}$ follow a ground truth joint distribution $\pi(\mathcal{P},\mathcal{Y})=\pi(\mathcal{Y}|\mathcal{P})\pi(\mathcal{P})$.
Problem Formulation. Let $h(\cdot)$ be a 3D semantic segmentation model that takes point cloud $\mathcal{P}$ as the input and outputs the class predictions $\hat{Y}=\{\hat{y}_{i}|i=0,1,...,N-1\}$ and the corresponding confidence scores $\hat{C}=\{\hat{c}_{i}|i=0,1,...,N-1\}$, i.e., $h(\mathcal{P})=(\hat{Y},\hat{C})$. Our goal is two-fold: (1) we would like to measure how well the 3D semantic segmentation model delivers the uncertainty estimates when making predictions, and (2) we anticipate obtaining a well-calibrated 3D semantic segmentation model that makes accurate predictions when the associated confidence scores are high. In theory, a perfect model calibration is defined as $\mathbb{P}(\hat{y}_{i}=y_{i}|\hat{c}_{i}=c)=c$, where $c\in[0,1]$ is the expected confidence value.
Objective. To better cater to real-world requirements, we resort to the non-probabilistic [^3] In the deep learning context, the non-probabilistic output $\mathbf{z}_{i}$ is often known as logits. output $\mathcal{Z}=\{\mathbf{z}_{i}|i=0,1,...,N-1\}$ from the 3D semantic segmentation model for calibration, without altering the model’s accuracy. The predicted probability $\hat{c}_{i}$ can be derived from $\mathbf{z}_{i}$ using a Softmax function, i.e., $\hat{c}_{i}=\sigma(\mathbf{z}_{i})$, with $\sigma(\cdot)$ denoting the Softmax operation. The overall objective is to produce a calibrated probability $\hat{v}_{i}$ for each point in $\mathcal{P}$, based on $\hat{y}_{i}$, $\hat{c}_{i}$, and $\mathbf{z}_{i}$.
@@ 3.1 Calibration Metrics
Expected Calibration Error (ECE). Guo et al. [36] proposed the ECE metric to assess the confidence calibration of a given neural network. Specifically, ECE measures the difference in expectation between confidence and accuracy:
@@
Based on the definition, a perfectly calibrated model has an ECE value of zero.
ECE for 3D Scene Understanding. In practice, we approximate the value of Eq. 1 by quantizing continuous variables into equally interval probability bins. Different from the conventional image classification task [36, 71], we treat each of the $N$ points [^4] It is worth noting that point clouds often contain an unequal number of points, i.e., take different $N$ values, due to acquisitions; we omit such a difference for simplicity. in $\mathcal{P}$ as unique samples. Assuming a total of $\tilde{N}$ point clouds in a dataset, we first calculate the ECE score of each point cloud and then average across all point clouds. Such a statistical binning takes the weighted average of the accuracy/confidence difference of each bin in each sample as follows:
@@ 3.2 Calibration Methods
Temperature Scaling (TempS). As has been widely verified in theory and practice, a simple extension of the Platt scaling [86, 78] is effective in improving the model calibration. Following [36], we use a single temperature parameter $T>0$ for all $S$ classes to re-scale the non-probabilistic output $\mathbf{z}_{i}$, that is:
@@
where $\sigma(\cdot)$ denotes the Softmax function. The $\max(\cdot)$ operation takes the maximum value at the $s$-th position over $S$ semantic classes. Conventionally, $T$ is learned by minimizing the log-likelihood loss on a validation set.
Logistic Scaling (LogiS). A more flexible version of the temperature scaling adopts more complex transformations during re-scaling. Guo et al. [36] proposed to use the logistic regression to adjust the non-probabilistic output $\mathbf{z}_{i}$, that is:
@@
where $\mathbf{W}$ and $\mathbf{b}$ are optimized based on the negative log-likelihood loss on a validation set. In this work, we adopt the vector scaling variant of logistic scaling, where the parameter $\mathbf{W}$ is restricted to a diagonal matrix.
Dirichlet Scaling (DiriS). Assuming the model’s outputs follow a Dirichlet distribution (rather than just single probability values), Kull et al. [61] further derived the Dirichlet scaling from logistic scaling, that is:
@@
Figure 2: Depth-correlated patterns in a ±$\pm 50$m LiDAR scene from SemanticKITTI [7]. (a) Ground truth. (b) Point-wise ECE scores. (c) Point-wise entropy scores.
@@
where $\mathbf{W}$ and $\mathbf{b}$ are parameters for a linear parameterization of the predicted probability $\sigma(\mathbf{z}_{i})$, and, similar to Eq. 3 and Eq. 4, $\mathbf{W}$ and $\mathbf{b}$ can be optimized based on the negative log-likelihood loss on a validation set.
@@
Meta-Calibration (MetaC). Ma et al. [71] proposed to combine a base calibrator (e.g., temperature scaling) with a bipartite ranking model for better calibration performance. Specifically, the prediction entropy calculated via $\hat{c}_{i}$ is used to select calibrators; the base calibrator will be used if the entropy is lower than a threshold, and, on the contrary, the predicted output will take random values. Theoretical analyses on high-probability bounds w.r.t. mis-coverage rate and coverage accuracy are presented in [71]. In practice, this is formulated as:
@@
$\hat{v}_{i}^{\text{MetaC}}=\left\{\begin{aligned} S^{-1}~{},&~{}~{}~{}\text{if}-c_{i}\log(c_{i})>\eta\\ \max_{s}\sigma(\frac{\mathbf{z}_{i}}{T})^{(s)}~{},&~{}~{}~{}\text{otherwise}\end{aligned}~{},\right.$ (6)
@@
where temperature $T>0$ is learned by minimizing the log-likelihood loss on a validation set. $\eta$ is a hand-crafted threshold for filtering high-entropy predictions. It is worth noting that meta-calibration, albeit proven effective in previous literature, will inevitably lose accuracy preservation. The first condition in Eq. 6 introduces randomness to model predictions, which is likely to be impractical regarding real-world, safety-critical applications, e.g., 3D scene understanding.
3.3 DeptS: Depth-Aware Scaling for Better 3D Calibration

### v2

@@ Abstract
Safety-critical 3D scene understanding tasks necessitate not only accurate but also confident predictions from 3D perception models. This study introduces Calib3D, a pioneering effort to benchmark and scrutinize the reliability of 3D scene understanding models from an uncertainty estimation viewpoint. We comprehensively evaluate 28 state-of-the-art models across 10 diverse 3D datasets, uncovering insightful phenomena that cope with both the aleatoric and epistemic uncertainties in 3D scene understanding. We discover that despite achieving impressive levels of accuracy, existing models frequently fail to provide reliable uncertainty estimates – a pitfall that critically undermines their applicability in safety-sensitive contexts. Through extensive analysis of key factors such as network capacity, LiDAR representations, rasterization resolutions, and 3D data augmentation techniques, we correlate these aspects directly with the model calibration efficacy. Furthermore, we introduce DeptS, a novel depth-aware scaling approach aimed at enhancing 3D model calibration. Extensive experiments across a wide range of configurations validate the superiority of our method. We hope this work could serve as a cornerstone for fostering reliable 3D scene understanding. Code and benchmark toolkit are publicly available [^1] https://github.com/ldkong1205/Calib3D.
Figure 1: Well-calibrated 3D scene understanding models are anticipated to deliver low uncertainties when predictions are accurate and high uncertainties when predictions are inaccurate. Existing 3D models [144] (UnCal) and prior calibration methods [38, 80] struggled to provide proper uncertainty estimates. Our proposed depth-aware scaling (DeptS) is capable of outputting accurate estimates, highlighting its potential for real-world usage. The plots shown are the point-wise expected calibration error (ECE) rates. The colormap goes from dark to light, denoting low and high error rates, respectively. Best viewed in colors.
 Lingdong and Xiang contributed equally to this work.
@@ 1 Introduction
The reliability of perception systems in real-world conditions is paramount. Safety-critical applications, such as autonomous driving and robot navigation, often rely on robust and accurate predictions from perception models [140, 107, 72, 64]. While learning-based perception models are widely adopted, they often struggle to provide reliable uncertainty estimates [36] and can exhibit over- or under-confidence [51]. This poor calibration fails to meet the demands of real-world applications [1, 88, 127], contradicting safety requirements in autonomous systems, where precise, confident predictions are critical for obstacle detection [106, 59, 126]. Similar concerns exist in safety-critical areas, e.g., surveillance [67, 86, 97], healthcare [83, 54, 82], and remote sensing [103, 35].
@@
Motivation. Despite these efforts, the reliability of 3D scene understanding models in providing uncertainty estimates [^2] In this work, for the sake of clarity, the terms uncertainty and confidence are used interchangeably, i.e., uncertainty $=1~{}-$ confidence. remains underexplored. 3D data, such as LiDAR and RGB-D camera inputs, are sparser and less structured than images [11, 2, 124]. Calib3D is designed to benchmark and study the reliability of 3D models through uncertainty estimation, focusing on both aleatoric and epistemic uncertainties to address real-world, safety-critical challenges. Specifically, our study emphasizes two key aspects:
@@
$\bullet$ Aleatoric Uncertainty in 3D. We examine how intrinsic factors, e.g., sensor noises [33, 59, 42] and point cloud density variations [114, 77, 16, 15], contribute to data uncertainty in 3D perception, which cannot be reduced by involving more data or using improved models. In Calib3D, we contribute a comprehensive study of 10 diverse 3D datasets, spanning different sensors, annotations, and scene settings, including driving, off-road, indoor, dynamic, synthetic/simulation, adverse weather conditions, etc.
@@
$\bullet$ Epistemic Uncertainty in 3D. Different from the rather unified network structures in 2D [44, 27], 3D scene understanding models encompass diverse structures due to the complex nature of 3D data processing. Our investigation in Calib3D extends to the model uncertainty associated with the diverse 3D architectures, highlighting the importance of addressing knowledge gaps in model training and data representation. A total of 28 state-of-the-art models are compared and analyzed, shedding light on the future development of more reliable 3D scene understanding models.
@@
Our analysis reveals that while 3D models often achieve high accuracy, their calibration falls short, a gap critical in safety-critical applications. While these models often achieve promising levels of accuracy, their calibration abilities – essential for trust in safety-critical applications – consistently fall short of the mark. As shown in Fig. 1, better calibration methods are needed to align model confidence with accuracy. Through a detailed examination of network capacity, LiDAR data representations, rasterization, and 3D data augmentation, we identify key areas for improvement.
@@
To further enhance uncertainty estimation capabilities, we propose a depth-aware scaling method called DeptS. Our method is motivated by the observation that uncalibrated models tend to have low accuracy in the middle-to-far region of the ego-vehicle, while, in the meantime, posing severe over-confident predictions. This problem, which is directly correlated with 3D scene structural information, inevitably leads to high calibration errors. To tackle this challenge, we design a depth-correlated temperature to dynamically adjust the logits distribution based on the depth information, exhibiting strong generalizability in calibrating 3D perception models. As demonstrated in Fig. 1, DeptS not only significantly improves calibration over uncalibrated models but also outperforms several existing methods [38, 66, 119, 80]. To encapsulate, this work is featured by the following seminar contributions:
[…]
@@ 2 Related Work
3D Scene Understanding. Holistic 3D perception underpins various real-world applications [11, 6]. Existing methods can be categorized based on 3D representations [115], including range view [84, 142, 18, 129, 61, 58, 133], bird’s eye view [141, 14, 143], sparse voxel [20, 144, 110, 48, 47, 19], and raw points [49, 112, 139, 98]. Recent work combines these representations [73, 131, 75] or fuses point clouds with other modalities (e.g., cameras, radars, IMU) [145, 53, 76, 94, 130] to enhance accuracy. While 3D perception has progressed on popular benchmarks, the reliability of these models in estimating uncertainty remains unexplored.
@@
Uncertainty Estimation. Quantifying uncertainties is crucial in real-world scenarios [4, 45], especially for safety-critical applications such as 3D scene understanding [91]. Methods for uncertainty estimation generally fall into various types, i.e., deterministic networks [81, 105], Bayesian methods [25, 113, 10, 46, 32], ensembles [68, 101, 39], and test-time augmentations [79, 5]. Recent studies pay attention to post-hoc approaches for calibrating uncertainties, which align with practical usages [36]. This work follows this line of research and extends efforts to 3D scene understanding, hoping to enlighten future works on this crucial topic.
@@
Network Calibration. As a prevailing research topic, numerous calibration methods have been proposed across various tasks, including image classification [38, 138, 66, 80, 89, 93, 65], semantic segmentation [26, 55, 119, 85, 9, 29], object detection [67, 86, 56, 90, 97], depth estimation [57, 52, 96, 117], remote sensing [103, 35], medical imaging [83, 54, 82], etc. However, the calibration of 3D scene understanding models, which relates closely to real-world applications, is rather overlooked in the literature. Dreissig et al. [30] made an initial study of SalsaNext [23] on the SemanticKITTI [7] dataset. To our knowledge, Calib3D is the first study of uncertainty estimation for 3D perception models, covering 28 state-of-the-art models across 10 datasets. We also propose DeptS, a novel depth-aware scaling method that effectively improves uncertainty calibration.
3D Robustness. The robustness of perception models has gained increasing attention, particularly in driving applications. Research has examined robustness to point cloud corruptions [59, 41, 40, 102, 12, 135, 125], depth corruptions [63, 34, 60], and multi-view images [126, 50, 17, 42]. Other works explore robustness against sensor failures [137, 37] and adversarial attacks [140, 128]. Unlike prior work, our focus is on 3D robustness from the perspective of uncertainty estimation. Calib3D establishes the first comprehensive benchmark in this area, aiming to guide future research in developing more reliable 3D perception models.
@@ 3 Calib3D
A typical point cloud data $\mathcal{P}=\{\mathbf{p}_{i},q_{i}|i=1,2,\dots,N\}$ contains $N$ points captured by the sensor, where $\mathbf{p}_{i}\in\mathbb{R}^{3}$ represents the Cartesian coordinates $(p^{x}_{i},p^{y}_{i},p^{z}_{i})$ and $q_{i}\in\mathbb{R}^{1}$ denotes the sensor reflection value, e.g., the laser intensity. For a learning-based system, the data is accompanied by semantic labels $\mathcal{Y}=\{y_{i}|i=1,2,\dots,N\}$ for each point in $\mathcal{P}$, with $y_{i}$ indicating one of $S$ pre-defined semantic classes. The random variables $\mathcal{P}$ and $\mathcal{Y}$ follow a ground truth joint distribution $\pi(\mathcal{P},\mathcal{Y})=\pi(\mathcal{Y}|\mathcal{P})\pi(\mathcal{P})$.
Problem Formulation. Let $h(\cdot)$ be a 3D model that takes a point cloud $\mathcal{P}$ as the input and outputs class predictions $\hat{Y}=\{\hat{y}_{i}|i=1,2,\dots,N\}$ along with confidence scores $\hat{C}=\{\hat{c}_{i}|i=1,2,\dots,N\}$, i.e., $h(\mathcal{P})=(\hat{Y},\hat{C})$. Our goal here is two-fold: (1) we want to measure how well the 3D model delivers the uncertainty estimates in its predictions; and (2) we anticipate obtaining a well-calibrated 3D perception model that aligns high confidence scores with accurate predictions. In theory, a perfect model calibration is defined as $\mathbb{P}(\hat{y}_{i}=y_{i}|\hat{c}_{i}=c)=c$, where $c\in[0,1]$ is the expected confidence value.
Objective. To better cater to the real-world requirement, we resort to the non-probabilistic [^3] In the deep learning context, the non-probabilistic output $\mathbf{z}_{i}$ is often known as logits. output $\mathcal{Z}=\{\mathbf{z}_{i}|i=1.2,\dots,N\}$ from the 3D semantic segmentation model for calibration, without altering the model’s accuracy. The predicted probability $\hat{c}_{i}$ can be derived from $\mathbf{z}_{i}$ using a Softmax function, i.e., $\hat{c}_{i}=\sigma(\mathbf{z}_{i})$, with $\sigma(\cdot)$ denoting the Softmax operation. The overall objective is to produce a calibrated probability $\hat{v}_{i}$ for each point in $\mathcal{P}$, based on $\hat{y}_{i}$, $\hat{c}_{i}$, and $\mathbf{z}_{i}$.
Figure 2: Depth-correlated patterns in a ±$\pm 50$m LiDAR-acquired scene from the SemanticKITTI [7] dataset. (a) Ground truth semantics. (b) Point-wise ECE scores. (c) Point-wise entropy scores.
@@ 3.1 Calibration Metrics
Expected Calibration Error (ECE). Guo et al. [38] introduced the ECE metric to assess the confidence calibration of a given neural network. Specifically, ECE measures the difference in expectation between confidence and accuracy:
@@
Based on the definition, a perfectly calibrated model will have an ECE value of zero.
ECE for 3D Scene Understanding. In practice, Eq. 1 is approximated by binning continuous variables into equally spaced probability intervals. Different from the conventional image classification task [38, 80], we treat each of the $N$ points [^4] Note that point clouds may contain varying numbers of points due to acquisitions; we omit such a difference for simplicity. in $\mathcal{P}$ as unique samples. Assuming a total of $\tilde{N}$ point clouds in a dataset, we first calculate the ECE score of each point cloud and then average across all point clouds. Such a statistical binning takes the weighted average of the accuracy/confidence difference of each bin as follows:
@@ 3.2 Calibration Methods
Temperature Scaling (TempS). As has been widely verified in theory and practice, a simple extension of the Platt scaling [95, 87] is effective in improving the model calibration. Following [38], a single temperature parameter $T>0$ is used to re-scale the non-probabilistic output $\mathbf{z}_{i}$:
@@
where $\sigma(\cdot)$ is the Softmax function, and $\max(\cdot)$ selects the maximum value over $S$ semantic classes. $T$ is learned by minimizing the log-likelihood loss on a validation set.
Logistic Scaling (LogiS). A more flexible version of the temperature scaling adopts more complex transformations during re-scaling. Guo et al. [38] proposed to use the logistic regression to adjust the non-probabilistic output $\mathbf{z}_{i}$:
@@
where $\mathbf{W}$ and $\mathbf{b}$ are optimized based on negative log-likelihood loss on a validation set. In this work, we adopt a vector scaling variant where $\mathbf{W}$ is diagonal.
Dirichlet Scaling (DiriS). Assuming the model’s outputs follow a Dirichlet distribution (rather than just single probability values), Kull et al. [66] further derived the Dirichlet scaling from logistic scaling, which is:
@@
where $\mathbf{W}$ and $\mathbf{b}$ are parameters for a linear parameterization of the predicted probability $\sigma(\mathbf{z}_{i})$, and similar to Eq. 3 and Eq. 4, $\mathbf{W}$ and $\mathbf{b}$ can be optimized based on the negative log-likelihood loss on a validation set.
@@
Meta-Calibration (MetaC). Ma et al. [80] combined a base calibrator (e.g., temperature scaling) with a bipartite ranking model for improved calibration. Specifically, prediction entropy is used to select calibrators; the base calibrator will be used if the entropy is lower than a threshold, and, on the contrary, the predicted output will take random values. Theoretical analyses on high-probability bounds w.r.t. mis-coverage rate and coverage accuracy are presented in [80]. In practice, this is formulated as:
@@
$\hat{v}_{i}^{\text{MetaC}}=\begin{cases}S^{-1},&\text{if }-c_{i}\log(c_{i})>\eta\\ \max_{s}\sigma(\frac{\mathbf{z}_{i}}{T})^{(s)},&\text{otherwise}\\ \end{cases}~{},$ (6)
@@
where temperature $T>0$ is learned via log-likelihood minimization on a validation set. $\eta$ is a hand-crafted threshold for filtering high-entropy predictions. It is worth noting that meta-calibration, albeit proven effective in previous literature, will inevitably lose accuracy preservation. The first condition in Eq. 6 introduces randomness to model predictions, which is likely to be impractical regarding real-world, safety-critical applications, e.g., 3D scene understanding.
3.3 DeptS: Depth-Aware Scaling for 3D Calibration